OpenStudy (anonymous):

help im stuck! e^8x=10 help me solve ;s

OpenStudy (anonymous):

Take natural log of both sides: \[\ln(e^{8x})=\ln(10) \implies 8x=\ln(10)\] Divide by 8 and you're done :P

OpenStudy (anonymous):

8x=ln10 X=ln10/8

OpenStudy (anonymous):

thank you!